# RT-DATR:Real-time Unsupervised Domain Adaptive Detection Transformer with Adversarial Feature Learning

### v1

@@
Abstract.
Despite domain-adaptive object detectors based on CNN and transformers have made significant progress in cross-domain detection tasks, it is regrettable that domain adaptation for real-time transformer-based detectors has not yet been explored. Directly applying existing domain adaptation algorithms has proven to be suboptimal. In this paper, we propose RT-DATR, a simple and efficient real-time domain adaptive detection transformer. Building on RT-DETR as our base detector, we first introduce a local object-level feature alignment module to significantly enhance the feature representation of domain invariance during object transfer. Additionally, we introduce a scene semantic feature alignment module designed to boost cross-domain detection performance by aligning scene semantic features. Finally, we introduced a domain query and decoupled it from the object query to further align the instance feature distribution within the decoder layer, reduce the domain gap, and maintain discriminative ability. Experimental results on various benchmarks demonstrate that our method outperforms current state-of-the-art approaches. Our code will be released soon.
real-time object detection, unsupervised domain adaptation, adversarial feature alignment
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandisbn: 978-1-4503-XXXX-X/2025/06submissionid: 2531ccs: Computing methodologies Object detection
Figure 1. Performance comparison of different benchmarks.
1. Introduction
@@
To tackle this challenge, an increasing body of research has focused on developing unsupervised domain adaptation (UDA) (Long et al., 2016) algorithms for object detection. These efforts include adversarial feature learning, image translation, and semi-supervised approaches based on the student-teacher framework. The primary goal of these methods is to leverage both labeled source domain data and unlabeled target domain data to train a detector that can generalize effectively in the target domain. For example, several influential studies integrate adversarial learning with object detection by introducing a domain classifier and a gradient reversal layer (GRL) (Ganin and Lempitsky, 2015) into the detector. This approach minimizes the classifier loss while maximizing the detector loss, thereby encouraging the learning of domain-invariant features and enhancing detection performance in the target domain (Chen et al., 2018). Additionally, methods based on the mean-teacher framework (Deng et al., 2021a) explore learning strategies through the interaction between teacher and student models, progressively adapting the object detector to cross-domain detection tasks.
Many existing domain adaptation methods are meticulously designed for CNN-based detectors (Tian et al., 2020; Ren et al., 2016). However, with the outstanding performance of transformers in object detection (Carion et al., 2020), transformer-based domain-adaptive object detection has gained significant attention and achieved remarkable progress. Unlike CNN-based detectors, transformer-based detectors address domain shifts by aligning feature distributions across various modules. This is accomplished through the integration of domain-specific queries or by leveraging text-to-image models to synthesize data resembling the target domain from the source domain (Huang et al., 2024). These approaches progressively reduce the domain gap, enabling superior performance in the target domain.
@@
To address these challenges, in this paper, we propose RT-DATR, which, to our knowledge, is the first real-time domain adaptive detection transformer. As shown in Figure 1, it achieved state-of-the-art performance on different benchmarks. We chose RT-DETR (Zhao et al., 2024) as our base detector because it is currently the most popular and widely used DETR-style real-time detector. The modules we introduced do not add any extra inference latency, ensuring the high performance of the model. Specifically, we propose a local object-level feature alignment (LOFA) module and a scene semantic feature alignment (SSFA) module. The former can achieve pixel-level local alignment based on multi-scale features while also enhancing the feature alignment of the object regions. The latter globally aligns the encoded high-level scene semantic features, enabling the learning of transferable domain-invariant features. We introduced a domain query and decoupled it from the object query, aiming to alleviate the domain gap through adversarial learning while maintaining discriminative ability during prediction. Additionally, we enhanced the model’s robustness by incorporating consistency loss in the Instance Feature Alignment (IFA) module. Our main contributions can be summarized as follows:
@@
  • We developed three distinct adversarial feature learning modules, which include components for aligning local object-level features and scene semantic features. These modules leverage adversarial learning both at the object level and globally to acquire domain-invariant features and enhance object transferability. Furthermore, we incorporate domain queries and consistency loss to further minimize domain gaps and enhance the model’s robustness. Notably, these modules are all infer-free.
@@
2. Related Work
2.1. Transformer-based Object Detection
DETR (Carion et al., 2020) is the first end-to-end object detector based on transformer. It achieves one-to-one set prediction by using the Hungarian matching algorithm without relying on components such as anchor boxes and non-maximum suppression. Although DETR (Carion et al., 2020) has the advantage of end-to-end prediction, it suffers from slow convergence and high computational costs. To address these issues, Deformable DETR (Zhu et al., 2020) introduces a deformable attention module that focuses attention around the reference point, significantly reducing computation cost and accelerating training convergence.
DINO (Zhang et al., 2022) employs a mixed query selection strategy for better query initialization and introduces a “look forward twice” scheme to refine box predictions. RT-DETR (Zhao et al., 2024) is the first transformer-based real-time detector that achieves significant speed improvements while maintaining high accuracy through efficient hybrid encoder design and minimal uncertainty query selection. RT-DETRv2 (Lv et al., 2024) introduces a series of “bag-of-freebies” to improve flexibility and practicality while optimizing training strategies to boost performance. RT-DETRv3 (Wang et al., 2024) proposes a hierarchical dense auxiliary supervision scheme to further enhance detection accuracy and convergence speed. In this paper, we selected RT-DETR as our base detector because it is currently the most popular DETR-style real-time detector. To ensure the efficiency of RT-DATR, we designed the modules to enhance cross-domain detection performance without introducing additional inference latency.
Figure 2. The architecture of RT-DATR. It consists of the base detector RT-DETR, along with three feature alignment modules: local object-level, scene semantic, and instance feature alignment modules. Object proposals are the object regions predicted by the model for object-level feature alignment.
2.2. UDA for Object Detection
@@ 2.2.1. UDA for CNN-based Detector
Unsupervised domain-adaptive object detection was first proposed in  (Chen et al., 2018), based on Faster RCNN (Ren et al., 2016) using two gradient reverse layer (GRL) (Ganin and Lempitsky, 2015) for optimization through minimum-maximum adversarial learning, achieving image-level and instance-level alignment. (Saito et al., 2019) proposed a strong-weak distribution alignment method, which focuses the adversarial alignment loss on images that are globally similar and puts less emphasis on aligning images that are globally dissimilar, achieving significant results. Considering that using domain classifiers to reduce the differences between the source and target domains can ignore the difficulties of these transferable features in processing classification and localization subtasks in object detection.Therefore, task-specific inconsistency alignment (Zhao and Wang, 2022) is proposed, by developing a new alignment mechanism in separate task spaces,improving the performance of the detector on both sub-tasks.
@@
Unlike adversarial feature learning, there are also some works that maximize the data information of the target domain by fully utilizing pseudo-label self-training. For example, PT (Chen et al., 2022) aims to capture the uncertainty of unlabeled target data from evolving teachers and guide students’ learning in a mutually beneficial manner, thereby avoiding the negative impact of pseudo box quality. CMT (Cao et al., 2023) proposed a framework that combines contrastive learning and mean teacher self-training, using pseudo labels to extract object level features and optimizing them through contrastive learning. Additionally, methods involving image transformation have been employed to reduce the domain gap. For instance, ConfMix(Mattolin et al., 2023) is the first approach to utilize a sample mixing strategy based on region-level detection confidence for adaptive object detector learning, adopting YOLOv5(Jocher et al., 2022) as its base detector.
2.2.2. UDA for Transformer-based Detector
@@
However, unlike CNN-based domain adaptive object detection, research on adopting DETR-style architectures for domain adaptive detection tasks remains relatively scarce. SFA(Wang et al., 2021) proposes a method for sequence feature alignment by introducing domain queries to align the features of the encoder and decoder, respectively, achieving promising results. O2Net(Gong et al., 2022) improves performance by mitigating the domain shift on the backbone and aligning the output features of the encoder. AQT (Huang et al., 2022) further integrates adversarial feature alignment methods into the detection transformer and implements space, channel, and instance-level feature alignment to produce domain-invariant features. BiADT(He et al., 2023) proposes a novel deformable attention and self-attention approach that aims to achieve bidirectional domain alignment and mutual information minimization by explicitly labeling domain-invariant and domain-specific features in image and object labeling sequences. DA-DETR(Zhang et al., 2023) achieves effective transformation from labeled source domain to unlabeled target domain by fusing CNN and Transformer information. MRT (Zhao et al., 2023) masks the multi-scale feature map of the target image through the encoder and auxiliary decoder of the student model, and reconstructs the features to help the student model capture target domain features and mitigate the impact of low-quality pseudo-bounding boxes. MTM (Weng and Yuan, 2024) proposes masked feature alignment methods to alleviate domain shift in a more robust way, which not only prevents training stagnation and leads to a robust pretrained model in the pretraining stage but also enhances the model’s target performance in the self-training stage.
@@
Some approaches adopt image-to-image translation methods, utilizing text-to-image generation models to produce more diverse data from the target domain, thereby enhancing the model’s generalization ability. Among them, Blenda (Huang et al., 2024) performs adaptive training by generating pseudo-samples from an intermediate domain along with their corresponding soft domain labels. Inspired by the above work, we explore for the first time the combination of adversarial feature learning strategies with a real-time transformer-based detector to enhance cross-domain detection performance.
3. Method
3.1. Task Defination and Overview
@@
In unsupervised domain adaptation, the training data includes labeled source images $D_{s}$, denoted as $\{x_{s},y_{s}\}$, and unlabeled target images $D_{t}$, denoted as $\{x_{t}\}$. Our goal is to train an object detector that performs well on target domain data. The overall architecture of our proposed method, RT-DATR, is illustrated in Figure 2. RT-DETR serves as our base detector, with three additional distinct feature alignment modules introduced for domain adaptation. In supervised learning, source domain images are fed into the model, and constraints are imposed by computing the loss between the predictions of RT-DETR and the ground truth, denoted as $\mathcal{L}_{det}$. For unsupervised learning, both source and target domain data are used as inputs, and optimization is performed through the LOFA, SSFA, and IFA modules to drive adversarial alignment of source and target features, denoted as $\mathcal{L}_{adv}$.
3.2. Local and Object-level Feature Alignment
@@
Adversarial feature alignment within different detector components has proven to be a simple yet effective approach for domain adaptation. However, studies reveal that directly aligning CNN backbone features in transformer-based detectors often degrades performance. (Wang et al., 2021). Subsequent research has thus shifted focus to aligning feature distributions in other modules or extracting domain-specific features through domain queries (Huang et al., 2022). In contrast, we emphasize the critical role of the backbone network as the primary module for learning domain-invariant features. The multi-scale features extracted by the backbone capture rich local information, such as textures and colors, which are essential for domain adaptation. To preserve this information, we adopt local feature alignment instead of global image-level alignment, ensuring robust feature matching. Furthermore, we enhance foreground object-level alignment, improving object transferability and boosting detection performance in cross-domain tasks.
@@
In Figure 2a, we illustrate how data from both the source and target domains are processed through a backbone network to extract multi-scale features denoted as ${f_{1},f_{2},f_{3}}$. Following this, a gradient reversal layer is integrated at each feature level, and a domain discriminator, which includes a $Conv$1x1 layer, is employed to derive the transformed multi-scale features ${f}^{\prime}_{1},{f}^{\prime}_{2},$ and ${f}^{\prime}_{3}$. Subsequently, domain classification is performed on each pixel, with the source domain labeled as 0 and the target domain as 1. The optimization objective for aligning local features is defined as follows:
(1) $\begin{split}\mathcal{L}_{loc}=\sum_{i=1}^{M}-dlog(D_{b}(f^{s(u,v)}_{i}))+\\ (1-d)log(1-D_{b}(f^{s(u,v)}_{i})))\end{split}$
@@
Where $d$ denotes the domain label, $D_{b}$ represents the domain discriminator of the backbone, $f_{i}$ stands for the $i$-th layer among the multi-scale features ($i=1,2,...,M$), and $s(u,v)$ indicates the positional information within the feature layer.
@@
(2) $\begin{split}\mathcal{L}_{obj}=\sum_{i=1}^{M}-w_{i}dlog(D_{b}(f^{s(u,v)}_{i}))+\\ w_{i}(1-d)log(1-D_{b}(f^{s(u,v)}_{i})))\end{split}$
@@
(3) $\mathcal{L}^{B}_{adv}=\mathcal{L}_{loc}+\mathcal{L}_{obj}$
3.3. Scene Semantic Feature Alignment
@@
(4) $\mathcal{L}^{E}_{adv}=\frac{1}{L}\sum_{j=0}^{L}(-dlog(D_{e}(p_{j}))+(1-d)log(D_{e}(p_{j})))$
Where $p_{j}$ represents the $j$-th layer feature (with $j=0,1,...,L$) output by the encoder, and $D_{e}$ denotes the discriminator within the encoder.
Table 1. Experimental results of the scenario normal weather to foggy weather: Cityscapes→Foggy Cityscapes. R34 and R50 represent ResNet-34 and ResNet-50, respectively.
@@
OADA (Yoo et al., 2022) R50 47.8 46.5 62.9 32.1 48.5 50.9 34.3 39.8 45.4
HT (Deng et al., 2023) R50 52.1 55.8 67.5 32.7 55.9 49.1 40.1 50.3 50.4
@@
SFA (Wang et al., 2021) R50 46.5 48.6 62.6 25.1 46.2 29.4 28.3 44.0 41.3
MTTrans (Yu et al., 2022) R50 47.7 49.9 65.2 25.8 45.9 33.8 32.6 46.5 43.4
DA-DETR (Zhang et al., 2023) R50 49.9 50.0 63.1 24.0 45.8 37.5 31.6 46.3 43.5
O2Net (Gong et al., 2022) R50 48.7 51.5 63.6 31.1 47.6 47.8 38.0 45.9 46.8
AQT (Huang et al., 2022) R50 49.3 52.3 64.4 27.7 53.7 46.5 36.0 46.4 47.1
MRT (Zhao et al., 2023) R50 52.8 51.7 68.7 35.9 58.1 54.5 41.0 47.1 51.2
MTM (Weng and Yuan, 2024) R50 51.0 53.4 67.2 37.2 54.4 41.6 38.4 47.7 48.9
@@
3.4. Instance Feature Alignment and Consistency Loss

### v2

# RT-DATR: Real-time Unsupervised Domain Adaptive Detection Transformer with Adversarial Feature Alignment

@@
Abstract
Despite domain-adaptive object detectors based on CNN and transformers have made significant progress in cross-domain detection tasks, it is regrettable that domain adaptation for real-time transformer-based detectors has not yet been explored. Directly applying existing domain adaptation algorithms has proven to be suboptimal. In this paper, we propose RT-DATR, a simple and efficient real-time domain adaptive detection transformer. Building on RT-DETR as our base detector, we first introduce a local object-level feature alignment module to significantly enhance the feature representation of domain invariance during object transfer. Additionally, we introduce a scene semantic feature alignment module designed to boost cross-domain detection performance by aligning scene semantic features. Finally, we introduced a domain query and decoupled it from the object query to further align the instance feature distribution within the decoder layer, reduce the domain gap, and maintain discriminative ability. Experimental results on various cross-domian benchmarks demonstrate that our method outperforms current state-of-the-art approaches. Code is available at https://github.com/Jeremy-lf/RT-DATR.
Figure 1: Performance comparison of different benchmarks.
1 Introduction
@@
To tackle this challenge, an increasing body of research has focused on developing unsupervised domain adaptation (UDA) [19] algorithms for object detection. These efforts include adversarial feature learning, image translation, and semi-supervised approaches based on the student-teacher framework. The primary goal of these methods is to leverage both labeled source domain data and unlabeled target domain data to train a detector that can generalize effectively in the target domain. For example, several influential studies integrate adversarial learning with object detection by introducing a domain classifier and a gradient reversal layer (GRL) [9] into the detector. This approach minimizes the classifier loss while maximizing the detector loss, thereby encouraging the learning of domain-invariant features and enhancing detection performance in the target domain [4]. Additionally, methods based on the mean-teacher framework [7] explore learning strategies through the interaction between teacher and student models, progressively adapting the object detector to cross-domain detection tasks.
Many existing domain adaptation methods are meticulously designed for CNN-based detectors [28, 23]. However, with the outstanding performance of transformers in object detection [2], transformer-based domain-adaptive object detection has gained significant attention and achieved remarkable progress. Unlike CNN-based detectors, transformer-based detectors address domain shifts by aligning feature distributions across various modules. This is accomplished through the integration of domain-specific queries or by leveraging text-to-image models to synthesize data resembling the target domain from the source domain [14]. These approaches progressively reduce the domain gap, enabling superior performance in the target domain.
@@
To address these challenges, in this paper, we propose RT-DATR, which, to our knowledge, is the first real-time domain adaptive detection transformer. As shown in Figure 1, it achieved state-of-the-art performance on different benchmarks. We chose RT-DETR [41] as our base detector because it is currently the most popular and widely used DETR-style real-time detector. The modules we introduced do not add any extra inference latency, ensuring the high performance of the model. Specifically, we propose a local object-level feature alignment (LOFA) module and a scene semantic feature alignment (SSFA) module. The former can achieve pixel-level local alignment based on multi-scale features while also enhancing the feature alignment of the object regions. The latter globally aligns the encoded high-level scene semantic features, enabling the learning of transferable domain-invariant features. We introduced a domain query and decoupled it from the object query, aiming to alleviate the domain gap through adversarial learning while maintaining discriminative ability during prediction. Additionally, we enhanced the model’s robustness by incorporating consistency loss in the Instance Feature Alignment (IFA) module. Our main contributions can be summarized as follows:
@@
  • We developed three distinct adversarial feature learning modules, which include components for aligning local object-level features and scene semantic features. These modules leverage adversarial learning both at the object level and globally to acquire domain-invariant features and enhance object transferability. Furthermore, we incorporate domain queries and consistency loss to further minimize domain gaps and enhance the model’s robustness. Notably, these modules are all inference-free.
@@
2 Related Work
2.1 Transformer-based Object Detection
DETR [2] is the first end-to-end object detector based on transformer. It achieves one-to-one set prediction by using the Hungarian matching algorithm without relying on components such as anchor boxes and non-maximum suppression. Although DETR [2] has the advantage of end-to-end prediction, it suffers from slow convergence and high computational costs. To address these issues, Deformable DETR [44] introduces a deformable attention module that focuses attention around the reference point, significantly reducing computation cost and accelerating training convergence.
DINO [38] employs a mixed query selection strategy for better query initialization and introduces a “look forward twice” scheme to refine box predictions. RT-DETR [41] is the first transformer-based real-time detector that achieves significant speed improvements while maintaining high accuracy through efficient hybrid encoder design and minimal uncertainty query selection. RT-DETRv2 [20] introduces a series of “bag-of-freebies” to improve flexibility and practicality while optimizing training strategies to boost performance. RT-DETRv3 [30] proposes a hierarchical dense auxiliary supervision scheme to further enhance detection accuracy and convergence speed. In this paper, we selected RT-DETR as our base detector because it is currently the most popular DETR-style real-time detector. To ensure the efficiency of RT-DATR, we designed the modules to enhance cross-domain detection performance without introducing additional inference latency.
2.2 UDA for Object Detection
@@
UDA for CNN-based Detector. Unsupervised domain-adaptive object detection was first proposed in  [4], based on Faster RCNN [23] using two gradient reverse layer [9] for optimization through minimum-maximum adversarial learning, achieving image-level and instance-level alignment. [24] proposed a strong-weak distribution alignment method, which focuses the adversarial alignment loss on images that are globally similar and puts less emphasis on aligning images that are globally dissimilar, achieving significant results. Considering that using domain classifiers to reduce the differences between the source and target domains can ignore the difficulties of these transferable features in processing classification and localization subtasks in object detection.Therefore, task-specific inconsistency alignment [40] is proposed, by developing a new alignment mechanism in separate task spaces,improving the performance of the detector on both sub-tasks.
@@
Unlike adversarial feature learning, there are also some works that maximize the data information of the target domain by fully utilizing pseudo-label self-training. For example, PT [3] aims to capture the uncertainty of unlabeled target data from evolving teachers and guide students’ learning in a mutually beneficial manner, thereby avoiding the negative impact of pseudo box quality. CMT [1] proposed a framework that combines contrastive learning and mean teacher self-training, using pseudo labels to extract object level features and optimizing them through contrastive learning. Additionally, methods involving image transformation have been employed to reduce the domain gap. For instance, ConfMix[21] is the first approach to utilize a sample mixing strategy based on region-level detection confidence for adaptive object detector learning, adopting YOLOv5[16] as its base detector.
@@
UDA for Transformer-based Detector. However, unlike CNN-based domain adaptive object detection, research on adopting DETR-style architectures for domain adaptive detection tasks remains relatively scarce. SFA[31] proposes a method for sequence feature alignment by introducing domain queries to align the features of the encoder and decoder, respectively, achieving promising results. O2Net[11] improves performance by mitigating the domain shift on the backbone and aligning the output features of the encoder. AQT [15] further integrates adversarial feature alignment methods into the detection transformer and implements space, channel, and instance-level feature alignment to produce domain-invariant features. BiADT[12] proposes a novel deformable attention and self-attention approach that aims to achieve bidirectional domain alignment. DA-DETR[39] achieves effective transformation from labeled source domain to unlabeled target domain by fusing CNN and Transformer information. MRT [42] masks the multi-scale feature map of the target image through the encoder and auxiliary decoder of the student model, and reconstructs the features to help the student model capture target domain features. MTM [32] proposes masked feature alignment methods to alleviate domain shift in a more robust way.
@@
Some approaches adopt image-to-image translation methods, utilizing text-to-image generation models to produce more diverse data from the target domain, thereby enhancing the model’s generalization ability. Among them, Blenda [14] performs adaptive training by generating pseudo-samples from an intermediate domain along with their corresponding soft domain labels. Inspired by the above work, we explore for the first time the combination of adversarial feature learning strategies with a real-time transformer-based detector to enhance cross-domain detection performance.
Figure 2: The architecture of RT-DATR. It consists of the base detector RT-DETR, along with three feature alignment modules: local object-level, scene semantic, and instance feature alignment modules. Object proposals are the object regions predicted by the model for object-level feature alignment.
3 Method
3.1 Task Defination and Overview
It is noteworthy that our focus is not on developing a new real-time object detector. Instead, our efforts are dedicated to improving the cross-domain detection performance of an existing transformer-based real-time detector, and this represents the first research endeavor in this field.
@@ 3.1 Task Defination and Overview
In unsupervised domain adaptation, the training data includes labeled source images $D_{s}$, denoted as $\{x_{s},y_{s}\}$, and unlabeled target images $D_{t}$, denoted as $\{x_{t}\}$. Our goal is to train an object detector that performs well on target domain data. The overall architecture of our proposed method, RT-DATR, is illustrated in Figure 2. RT-DETR serves as our base detector, with three additional distinct feature alignment modules introduced for domain adaptation. In supervised learning, source domain images are fed into the model, and constraints are imposed by computing the loss between the predictions of RT-DETR and the ground truth, denoted as $\mathcal{L}_{det}$. For unsupervised learning, both source and target domain data are used as inputs, and optimization is performed through the LOFA, SSFA, and IFA modules to drive adversarial fetaure alignment, denoted as $\mathcal{L}_{adv}$.
3.2 Local and Object-level Feature Alignment
@@
Adversarial feature alignment within different detector components has proven to be a simple yet effective approach for domain adaptation. However, studies reveal that directly aligning CNN backbone features in transformer-based detectors often degrades performance [31]. Subsequent research has thus shifted focus to aligning feature distributions in other modules or extracting domain-specific features through domain queries [15]. In contrast, we emphasize the critical role of the backbone network as the primary module for learning domain-invariant features. The multi-scale features extracted by the backbone capture rich local information, such as textures and colors, which are essential for domain adaptation. To preserve this information, we adopt local feature alignment instead of global image-level alignment, ensuring robust feature matching. Furthermore, we enhance foreground object-level alignment, improving object transferability and boosting detection performance in cross-domain tasks.
@@
In Figure 2a, we illustrate how data from both the source and target domains are processed through a backbone network to extract multi-scale features denoted as $f_{1},f_{2},f_{3}$}. Following this, a gradient reversal layer is integrated at each feature level, and a domain discriminator, which includes a ConvConv1x1 layer, is employed to derive the transformed multi-scale features ${f}^{\prime}_{1},{f}^{\prime}_{2},$ and ${f}^{\prime}_{3}$. Subsequently, domain classification is performed on each pixel, with the source domain labeled as 0 and the target domain as 1. The optimization objective for aligning local features is defined as follows:
$\begin{split}\mathcal{L}_{loc}=\sum_{i=1}^{M}-dlog(D_{b}(f^{s(u,v)}_{i}))+\\ (1-d)log(1-D_{b}(f^{s(u,v)}_{i}))\end{split}$ (1)
@@
Where $d$ denotes the domain label, $D_{b}$ represents the domain discriminator of the backbone, $f_{i}$ stands for the $i$-th layer among the multi-scale features (i=1,2,,Mi=1,2,...,M), and $s(u,v)$ indicates the positional information within the feature layer.
@@
$\begin{split}\mathcal{L}_{obj}=\sum_{i=1}^{M}-w_{i}dlog(D_{b}(f^{s(u,v)}_{i}))+\\ w_{i}(1-d)log(1-D_{b}(f^{s(u,v)}_{i}))\end{split}$ (2)
@@
$\mathcal{L}^{B}_{adv}=\mathcal{L}_{loc}+\mathcal{L}_{obj}$ (3)
3.3 Scene Semantic Feature Alignment
@@
$\mathcal{L}^{E}_{adv}=\frac{1}{L}\sum_{j=0}^{L}(-dlog(D_{e}(p_{j}))+(1-d)log(1-D_{e}(p_{j})))$ (4)
Where $p_{j}$ represents the $j$-th layer feature (with j=0,1,,Lj=0,1,...,L) output by the encoder, and $D_{e}$ denotes the discriminator within the encoder.
@@
OADA [35] R50 47.8 46.5 62.9 32.1 48.5 50.9 34.3 39.8 45.4
HT [8] R50 52.1 55.8 67.5 32.7 55.9 49.1 40.1 50.3 50.4
Source Only (YOLOv5) CSP-D53 34.8 37.6 48.7 14.3 30.1 8.8 14.6 28.1 27.1
ConfMix [21] CSP-D53 45.0 43.4 63.6 27.3 45.8 40.0 28.6 33.5 40.8
@@
SFA [31] R50 46.5 48.6 62.6 25.1 46.2 29.4 28.3 44.0 41.3
MTTrans [37] R50 47.7 49.9 65.2 25.8 45.9 33.8 32.6 46.5 43.4
DA-DETR [39] R50 49.9 50.0 63.1 24.0 45.8 37.5 31.6 46.3 43.5
O2Net [11] R50 48.7 51.5 63.6 31.1 47.6 47.8 38.0 45.9 46.8
AQT [15] R50 49.3 52.3 64.4 27.7 53.7 46.5 36.0 46.4 47.1
MRT [42] R50 52.8 51.7 68.7 35.9 58.1 54.5 41.0 47.1 51.2
MTM [32] R50 51.0 53.4 67.2 37.2 54.4 41.6 38.4 47.7 48.9
@@
Table 1: Experimental results of the scenario normal weather to foggy weather: Cityscapes→Foggy Cityscapes. R34, R50 and CSP-D53 represent ResNet-34, ResNet-50 and CSP-Darknet53 respectively.
3.4 Instance Feature Alignment and Consistency Loss